# SLayR: Scene Layout Generation with Rectified Flow

### v1

@@ Appendix B Implementation of Partial Conditioning
Finally, we note that this algorithm is a generalization of the inference algorithm presented in  Sec. 3.3 of our main paper, which is the special case when no conditioning is provided ($\mathbf{y}_{i}\equiv\varnothing$)
0:  Input parameters description
0:  Output description
1:  conditionedInference( $P_{i}$ , $\mathbf{y}_{i}$ ):
@@
13:  end while
14:  Return $\mathbf{x}_{i}(1)$
@@
Appendix C Training Data and Hyperparameters
@@
Appendix D Ablation of Image Generators
To control for the effect of the conditional generation component on our evaluation, we also run our text-to-image pipeline on competitors to InstanceDiffusion  [45], namely LMD+  [21], GliGEN  [20], and BoxDiff  [48], using the same image layouts we used in our main paper.
We show the generated image metrics for LMD+ in  Tab. 2. Like in our main paper, these metrics are inconclusive, and furthermore, the rankings of different layout generation methods differ depending on the downstream image generator.
LMD+
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$)
LayoutTransformer 0.730.730.730.730.73 $1.26$ $0.69$DeclareFontSeriesDefault[rm]bfb7.81
GPT4o $1.73$ 1.171.171.171.171.17 $0.69$ 7.667.667.667.667.66
Fixed Layout $3.81$ $1.93$ $0.77$ $7.06$
No Layout $1.82$ $1.87$ $0.83$ $7.33$
@@ Appendix D Ablation of Image Generators
GliGEN
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$)
[…]
GPT4o $3.49$ $1.30$ $0.63$ 8.138.138.138.138.13
Fixed Layout $2.88$ $2.51$ $0.68$ $6.57$
No Layout $1.85$ $1.89$ $0.82$ $7.51$
[…]
BoxDiff
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$)
LayoutTransformer $1.37$ 1.621.621.621.621.62 $0.53$DeclareFontSeriesDefault[rm]bfb9.049.049.049.04
GPT4o 1.231.231.231.231.23 $1.77$ $0.63$ $7.40$
Fixed Layout $2.32$DeclareFontSeriesDefault[rm]bfb1.421.421.421.42 0.520.520.520.520.52 $6.92$
No Layout $1.82$ $1.87$ $0.83$ $7.33$
[…]
Table 2: Metrics evaluated on LMD+, GliGEN, and BoxDiff. For LMD+, our method is lowest in FID and KID. Different from the results on the InstanceDiffusion-generated images shown in Tab. 1, it is now also lowest in CMMD. For GliGEN, our method now performs second-best on FID and KID, but best on CMMD and IS. This is opposite to the results of InstanceDiffusion-generated images shown in Tab. 1 For BoxDiff, our method is lowest in FID, we are not the lowest in KID anymore. We now perform the best in CMMD and second best in IS, unlike in images generated from InstanceDiffusion.
We also run a human evaluation on the images from each of these three generators, shown in  Fig. 10. Although our results are not identical between image generation methods, there are several important commonalities across all four of these methods: our method is significantly more plausible than LayoutTransformer, and has significantly more variety than GPT4o and No Layout. We even outperform GPT4o in terms of plausibility for LMD+ and GliGEN. These results support that our method offers an effective balance between plausibility and variety, regardless of the choice of image generator. These also show that our human evaluation procedure can help control for the choice of image generator.
Refer to caption Refer to caption Refer to caption
Figure 10: Human evaluations of plausibility and variability of our pipeline using the LMD+, GliGEN, and BoxDiff generators. Across these evaluators, we see that our method has higher plausibility than LayoutTransformer, and higher variety than GPT4o and No Layout. In the case of LMD+ and GliGEN, we even have higher plausibility than GPT4o
Appendix E LayoutTransformer Temperature
Throughout our main paper, we adapted LayoutTransformer as faithfully as possible to the original work, meaning that we kept the temperature parameter equal to one. However, the question arises whether the generated layouts would be higher quality at lower temperatures, where the model’s output is more stable. Therefore, we examined the model’s performance when setting the temperature to the lowest possible setting of zero.
The results in  Tab. 3 show that even though layout object positions improve at the lowest temperature, although this comes at the cost of poorer object numeracy and a complete lack of variance among produced layouts for the same prompt. The changes in the robust traditional metrics were also inconclusive, with FID worsening and KID improving at the lowest temperature. Furthermore, LayoutTransformer with a temperature of zero still performed worse than our method on FID, KID, and our layout measurement metrics. In  Fig. 11, we show qualtatively that the final images produced at lower temperatures are still lacking in plausibility.
Even in this best-case scenario when we select the lowest temperature for optimal stability, we are still not measuring a decisive improvement across numerical metrics. Therefore, we saw no clear indication that setting the temperature parameter lower than the default of one would lead to missed improvements in model performance on the human evaluation, and kept the temperature at its original setting of one to remain as faithful as possible to the prior work.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(10^{-11})(\uparrow)$ $l_{\text{Pos}}^{(2)}(10^{-11})(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
LayoutTransformer temp= 1 0.440.440.440.440.44 $0.94$DeclareFontSeriesDefault[rm]bfb1.341.341.341.34 $8.34$DeclareFontSeriesDefault[rm]bfb0.900.900.900.90 $3.09$ $1.21$DeclareFontSeriesDefault[rm]bfb231231231231
LayoutTransformer temp= 0 $0.48$ 0.920.920.920.920.92 $1.77$DeclareFontSeriesDefault[rm]bfb8.748.748.748.74 $4.11$ 3.733.733.733.733.73 1.531.531.531.531.53 00
[…]
Table 3: Comparison of metrics LayoutTransformer with a temperature of one (model default) and a temperature of zero. Even when the temperature is zero, we see that our method still performs better than the LayoutTransformer on the FID and KID metrics, which were shown to be robust in  Tab. 1. We also perform better than temp = $1$ in $O_{\text{Num}}$, $l_{\text{Pos}}^{(1)}$, $l_{\text{Pos}}^{(2)}$, and $\sigma^{2}_{\text{Pos}}$.
@@ Appendix F GPT4o Temperature
We experimented with increasing the temperature from $0.2$ in increments of $0.1$. We found that at a temperature of one, GPT4o failed to produce a parsable layout $14\%$ of the time. However, these mistakes were easy to catch and query the model again. Temperatures higher than $1$ caused more frequent parsing failures, and began to produce long, tangential sentences rather than proper object labels. It was not possible to heuristically filter these responses, and so we settled on a temperature of one as a reasonable upper limit for operation temperature of GPT4o on this task.
@@
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(10^{-11})(\uparrow)$ $l_{\text{Pos}}^{(2)}(10^{-11})(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
GPT4o temp=$0.25$ 0.940.940.940.940.94 $0.99$DeclareFontSeriesDefault[rm]bfb1.341.341.341.34 7.877.877.877.877.87 $3.71$DeclareFontSeriesDefault[rm]bfb4.374.374.374.37 1.491.491.491.491.49 $93$
GPT4o temp=1 $1.47$ $1.62$ $1.35$DeclareFontSeriesDefault[rm]bfb7.967.967.967.96 2.862.862.862.862.86 $4.02$ $1.35$ 142142142142142
@@
Table 4: Comparison of metrics GPT4o with a temperature of $0.25$ (adapted model default) and one (highest stable temperature). At increased temperatures, GPT4o performs worse on the FID and KID metrics, which were shown to be robust in  Tab. 1. Although increasing the temperature of GPT4o improves $O_{\text{Num}}$ (the object frequencies are closer to the ground truth) and $\sigma^{2}_{\text{Pos}}$ (the layouts are more varied overall), performance on $l_{\text{Pos}}^{(1)}$ and $l_{\text{Pos}}^{(2)}$ drops (the positions of the objects are less plausible). Our method still performs better in $O_{\text{Num}}$, $l_{\text{Pos}}^{(1)}$, $l_{\text{Pos}}^{(2)}$, and $\sigma^{2}_{\text{Pos}}$ than GPT4o with a raised temperature.
@@ Appendix H Comparison to DDIM
We provide an example here, from a model with an identical architecture to our presented model (including all hyperparameters specified in  Appendix C, except it is trained with a DDIM training objective and performs DDIM inference (with a log-linear noise schedule from $\sigma=0.02$ to $\sigma=1$). This is not an exhaustive search by any means, but is intended as a point-of-reference for other researchers.
We show our statistics in  Tab. 6, and some visual examples from the model in  Fig. 14. This model performs worse than ours for numerical metrics, which were robust against adversarial examples in  Sec. 5.5.
We are not entirely sure why we observe this performance between the two techniques. We speculate that the straighter transit paths of samples rectified flow  [24] increases the model’s ability to effectively learn high dimensional data like the PCA-reduced CLIP embeddings. However, this will require more investigation to confirm.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(10^{-11})(\uparrow)$ $l_{\text{Pos}}^{(2)}(10^{-11})(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
Ours (DDIM) $0.95$ $8.60$ $1.77$DeclareFontSeriesDefault[rm]bfb7.897.897.897.89 $4.33$ $0.25$ $0.01$DeclareFontSeriesDefault[rm]bfb239239239239
@@
Table 6: Generated image metrics, and our generated layout numerical metrics applied on our model architecture with DDIM or rectified flow. Our model performs better on image metrics that were robust to adversarial examples in  Tab. 1, i.e., FID and KID. DDIM performs poorly on $O_{\text{Num}}$, corroborating our observations of improper object labels in  Fig. 14, which also is responsible for the poor scores in $l_{\text{Pos}}^{(1)}$ and $l_{\text{Pos}}^{(2)}$. Although $\sigma^{2}_{\text{Pos}}$ is higher for DDIM, this is at the cost of the layouts being largely nonsense (see  Fig. 14)
@@
Appendix I Additional Images and Layouts
@@
Appendix J Print-ready Main Results Diagram

### v2

@@ 9 Implementation of Partial Conditioning
We construct the drift vector $\mathbf{d}_{i}$ which encodes the directional constraints. We begin by initializing $\mathbf{d}_{i}$ to 00 in all entries. Then, we add constraints. For example, if there is a constraint that bounding box $j$ must be left of bounding box $j^{\prime}$, then
$\mathbf{d}_{i}^{j}\leftarrow\mathbf{d}_{i}^{j}+(\lambda,0,0,0\mathbin{\|}0\mathbin{\|}0)\,,$ (15)
$\mathbf{d}_{i}^{j^{\prime}}\leftarrow\mathbf{d}_{i}^{j^{\prime}}+(-\lambda,0,0,0\mathbin{\|}0\mathbin{\|}0)\,,$ (16)
where $\lambda$ is a small constant.
In the special case when no conditioning is provided or directional constraints are provided ($\mathbf{y}_{i}\equiv\varnothing,\mathbf{d}_{i}=0$), this algorithm is identical to the rectified flow inference presented in  Sec. 3.3 of our main paper.
1:  conditionedInference( $P_{i}$ , $\mathbf{y}_{i}$ , $\mathbf{d}_{i}$ ):
@@
13:     $\mathbf{x}_{i}(t)\leftarrow\mathbf{x}_{i}(t)+\mathbf{d}_{i}$ //Apply drift for all given directional constraints
14:  end while
15:  Return $\mathbf{x}_{i}(1)$
@@
10 Training Data and Hyperparameters
@@
11 LayoutTransformer Temperature
Throughout our main paper, we maintained LayoutTransformer defaulttemperature parameter equal to one. However, the question arises whether the generated layouts would be higher quality at lower temperatures, where the model’s output is more stable. As shown in  Tab. 3 even when we select the lowest temperature of zero for optimal stability, we are still not measuring a decisive improvement across numerical metrics, therefore we kept the temperature at its original setting of one to remain as faithful as possible to the prior work.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(10^{-11})(\uparrow)$ $l_{\text{Pos}}^{(2)}(10^{-11})(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
LayoutTransformer temp= 1 0.440.440.440.440.44 $0.94$ $1.34$DeclareFontSeriesDefault[rm]bfb0.900.900.900.90 $3.09$ $1.21$DeclareFontSeriesDefault[rm]bfb231231231231
LayoutTransformer temp= 0 $0.48$ 0.920.920.920.920.92 $1.77$ $4.11$ 3.733.733.733.733.73 1.531.531.531.531.53 00
@@ 11 LayoutTransformer Temperature
Table 3: Comparison of metrics LayoutTransformer with a temperature of one (model default) and a temperature of zero. Even when the temperature is zero, we see that our method still performs better across our metrics.
@@ 12 GPT4o Temperature
We experimented with increasing the temperature from $0.2$ in increments of $0.1$. We found that at a temperature of $1$, GPT4o failed to produce a parsable layout $14\%$ of the time. However, these mistakes were easy to catch and query the model again. Temperatures higher than $1$ caused more frequent parsing failures, and began to produce long, tangential sentences rather than proper object labels. Without a method to heuristically filter these responses, we settled on a temperature of one as a reasonable upper limit for operation temperature of GPT4o on this task.
@@
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(10^{-11})(\uparrow)$ $l_{\text{Pos}}^{(2)}(10^{-11})(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
GPT4o temp=$0.25$ 0.940.940.940.940.94 0.990.990.990.990.99 1.341.341.341.341.34 $3.71$ 4.374.374.374.374.37 1.491.491.491.491.49 $93$
GPT4o temp=1 $1.47$ $1.62$ $1.35$ 2.862.862.862.862.86 $4.02$ $1.35$ 142142142142142
@@
Table 4: Comparison of metrics GPT4o with a temperature of $0.25$ (adapted model default) and one (highest stable temperature). At increased temperatures, GPT4o performs worse on the FID and KID metrics. Although increasing the temperature of GPT4o improves $O_{\text{Num}}$ (the object frequencies are closer to the ground truth) and $\sigma^{2}_{\text{Pos}}$ (the layouts are more varied overall), performance on $l_{\text{Pos}}^{(1)}$ and $l_{\text{Pos}}^{(2)}$ drops (the positions of the objects are less plausible). Our method still performs better in all displayed metrics.
@@ 14 Comparison to DDIM
We provide an example here, from a model with an identical architecture to our presented model (including all hyperparameters specified in  Sec. 10, except it is trained with a DDIM training objective and performs DDIM inference (with a log-linear noise schedule from $\sigma=0.02$ to $\sigma=1$). This is not an exhaustive search by any means, but is intended as a point-of-reference for other researchers.
We show our statistics in  Tab. 6, and some visual examples from the model in  Fig. 14. We speculate that the straighter transit paths of samples rectified flow  [27] increases the model’s ability to effectively learn high dimensional data like the PCA-reduced CLIP embeddings.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(\uparrow)$ $l_{\text{Pos}}^{(2)}(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
Ours (DDIM) $0.95$ $8.60$ $1.77$ $7.89$ $4.33$ $0.01$DeclareFontSeriesDefault[rm]bfb239239239239
@@
Table 6: Generated image metrics, and our generated layout numerical metrics applied on our model architecture with DDIM or rectified flow. Our model performs better on everything except positional variance $\sigma^{2}_{\text{Pos}}$, but this is at the cost of the layouts being largely nonsense (see  Fig. 14)
@@
15 Additional Images and Layouts
@@
16 Print-ready Main Results Diagram